OpenStudy (anonymous):

Find the values of x and y that make these triangles congruent by HL. A. x = 2, y = 2 B. x = 4, y = 4 C. x = 5, y = 5 D. x = 3, y = 3 Will medal and fan if correct!!!! :)

OpenStudy (danjs):

no diagram, but set the corresponding legs or angles equal to form some equations to solve

OpenStudy (danjs):

for 2 variables, need 2 equations

OpenStudy (anonymous):

Sorry there is a diagram, I forgot to add it. One sec

undeadknight26 (undeadknight26):

Seems like you got help :D Goodluck!

OpenStudy (danjs):

ok, what does "by HL " mean ?

OpenStudy (anonymous):

By Hypotenuse-Leg Theorem

OpenStudy (danjs):

oh, ok Two triangles ABC and AED the corresponding parts are congruent (the same)

OpenStudy (anonymous):

Right

OpenStudy (danjs):

They already give you that BC = ED, those two dashes mean same

OpenStudy (danjs):

That is first equation BC = ED y = x

OpenStudy (anonymous):

My second equation is 2x + 3 = 3y + 1

OpenStudy (danjs):

right, y = x and 2x + 3 = 3y +1

OpenStudy (anonymous):

But I don't know what to do next. :(

OpenStudy (danjs):

the first equation tells you that y=x, you can replace y in the second by y=x, x 2x + 3 = 3y + 1 2x + 3 = 3x + 1

OpenStudy (danjs):

If y = (something) You can replace y with all the (something) anywhere in the other equation, to reduce to one variable

OpenStudy (danjs):

you good? on solving that now?

OpenStudy (anonymous):

I think the answer is either A or B. When checked on the calc they both answers for x and y are the same.

OpenStudy (danjs):

yes they have to be, remember the first equation , y = x

OpenStudy (anonymous):

Ik, I mean A, the equations both equal 7, but for B. they both equal 13, but C and D are different answers. Thus leaving A and B.

OpenStudy (danjs):

the first equation tells you that y=x, you can replace y in the second by y=x 2x + 3 = 3y + 1 2x + 3 = 3x + 1 you can now solve that for x, after replaceing y by y=x 2x + 3 = 3x + 1

OpenStudy (danjs):

3x - 2x = 3 - 1 x = 2

OpenStudy (anonymous):

I got it! Thanks that helped a lot!!

OpenStudy (danjs):

welcome, if they dont give you y = x, but y = 5x + 3,. just put the 5x+3 in for y and solve, Substitution then use that value in either original equation to calculate the other variable